OpenStudy (anonymous):

Integral of e^-2x cosx

OpenStudy (anonymous):

I sense parts

OpenStudy (anonymous):

\[e ^{-2x} \cos x\]

OpenStudy (anonymous):

is there any way to do substitution?

OpenStudy (anonymous):

Because I don't think we've learnt parts yet

OpenStudy (anonymous):

You don't think? lol

OpenStudy (anonymous):

the integral of \[e ^{-2x} \cos x\] is\[\frac{ e ^{-2x} }{ -2 }\] right?

OpenStudy (anonymous):